OpenStudy (anonymous):

if t=pi/4 what is cos and sin

OpenStudy (anonymous):

sin(1/4pi) = cos(1/4pi) = \[{ 1 \over 2} \sqrt{2}\]

OpenStudy (anonymous):

\[\cos(\pi/4)\ \approx .99990605\] \[\sin(\pi/4)\approx .01370735\]

OpenStudy (anonymous):

remember to put your calculator on radians

OpenStudy (anonymous):

That's true, he did not specify one way or the other though.

OpenStudy (anonymous):

fair enough, but when you see pi it's usually radians :)

OpenStudy (anonymous):

ok i have a similar question t=pi/2, what is cos, and sin...why would sin be cos(1/2pi)

OpenStudy (anonymous):

look at the link I posted , the answer is there.

OpenStudy (anonymous):

Yes, he is right. The sin value is the y value, and cos in the x value, using your unit circle, you can find the points needed.

OpenStudy (anonymous):

fiddle, the link says pi/2 (0,1)...how can i interpet that with relevance to my question

OpenStudy (anonymous):

oh now that makes sense thx